### v1

@@ Appendix A Lagrange multiplier stabilisation for a conforming approximation
We report on experiments performed using the $\boldsymbol{\mathrm{H}}$-conforming stabilised scheme with Lagrange multipliers proposed in [41]. We test different stabilised schemes based on $\mathbb{P}_{0}$ and discontinuous $\mathbb{P}_{1}$ for the Lagrange multiplier.
@@
Table A.1: Example A. Error history for a finite element family with $\mathbb{P}^{2}_{2}-\mathbb{P}_{1}-\mathbb{P}_{0}-\mathbb{P}_{2}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})-$norm.
In the conforming case, the discrete inf-sup condition to be satisfied is given by
However, in [41] it is shown that despite choosing stable inf-sup elements such as Taylor–Hood, Mini-element, $\mathbb{P}_{2}-\mathbb{P}_{0}$, etc, together with a typical choice for the Lagrange multiplier space as above, this condition may not be satisfied. To circumvent this difficulty, one can either enrich the velocity space with bubbles having compact support along $\Gamma_{\mathrm{m}}$ (see [41] for details), or add suitable residual stabilisation in the discrete problem (see, for example [42, 40]). We adopt the latter option. We define generic finite element spaces $\boldsymbol{\mathrm{H}}_{h}\subset\boldsymbol{\mathrm{H}}$, $Q_{h}\subset Q$, $W_{h}\subset W$ and $Z_{h}\subset Z$ for the velocity, pressure, Lagrange multiplier, and concentration, respectively. Following [40], we first define the following mesh-dependent bilinear form $d_{h}:W_{h}\times W_{h}\to\mathbb{R}$:
$d_{h}(\lambda_{h},\xi_{h}):=\sum_{e\in\mathcal{E}_{\mathrm{m}}}h_{e}\int_{e}\lambda_{h}\,\xi_{h}\,\mathrm{d}s,\qquad\forall\lambda_{h},\xi_{h}\in W_{h}.$ (46)
$a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+\widetilde{a}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b(\boldsymbol{v}_{h},(p_{h},\lambda_{h}))+s_{1}((\boldsymbol{u}_{h},p_{h},\lambda_{h}),\boldsymbol{v}_{h})$ $=0,$ (47)
Note that for the conforming method, the discrete quantities $\xi_{h},\,\boldsymbol{\sigma}_{h}\boldsymbol{n}\cdot\boldsymbol{n},\,\nabla\boldsymbol{v}_{h}\boldsymbol{n}\cdot\boldsymbol{n},\,q_{h}\mathbb{I}\boldsymbol{n}\cdot\boldsymbol{n}$ all belong to $W_{h}$. Note also that, for a Navier–Stokes model with slip boundary condition, [42] proved that choosing $\delta=0$ and $\alpha_{0}$ lower than a threshold yields a stable method. As in [23], for $\delta=1$ we have symmetry, however a smallness condition on $\alpha_{0}$ is needed for sake of stability. For $\delta=-1$ we have the anti-symmetric variation of the method [16, 4], whose main advantage is the unconditional stability with respect to $\alpha_{0}$.
First let us consider the same domain and exact solution as in Test 7.1 and study the convergence of the conforming scheme using Taylor–Hood elements together with piecewise linear or constant discontinuous elements for the Lagrange multiplier. We also consider stabilised and non-stabilised formulations in order to test the robustness of the scheme.
@@
88 0.707 7.29e-01 $\star$ 5.21e-01 $\star$ 1.$0$star 8.25e-02 $\star$ 6
@@
Table A.2: Example A. Error history for a finite element family with $\mathbb{P}^{2}_{2}-\mathbb{P}_{1}-\mathbb{P}_{1}-\mathbb{P}_{2}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})-$norm.
The numerical results portrayed in Tables A.1A.2 clearly confirm the theoretical $O(h^{k+1})$-convergence in the energy norm similarly to the observed/predicted in [42, 40]. The blocks in Table A.1 show the error history displaying number of degrees of freedom, individual absolute errors, rates of convergence, and Newton iteration counts for the conforming discretisation using Taylor–Hood approximation of velocity-pressure, together with piecewise discontinuous elements for the Lagrange multiplier (on a submesh conforming with the bulk mesh), and piecewise quadratic and continuous functions for the concentration. However, the choice of $\mathbb{P}_{0}$ for the Lagrange multiplier shows an experimental rate of $\mathcal{O}(h^{1.5})$ for all the cases.
On the other hand, the results of using linear discontinuous elements, presented in Table A.2 show a noticeable deterioration of the convergence for the Lagrange multiplier when the stabilisation is removed. In turn, an optimal rate of convergence $O(h^{k+1})$ is achieved with stabilisation.
Appendix B Divergence-free non-conforming test
Finally, we use the lowest-order Crouzeix–Raviart elements ($\mathbb{CR}_{1}-\mathbb{P}_{0}$ for velocity-pressure pairs), characterised by velocities being nonconforming with $\boldsymbol{\mathrm{H}}^{1}(\Omega)$ instead of the div-free pair $\mathbb{BDM}_{1}-\mathbb{P}_{0}$. In addition, we use piecewise constants to approximate the Lagrange multiplier. Table B.1 describes the behaviour of the scheme with a stabilisation parameter $\alpha_{0}=20$, indicating a similar accuracy as in the $\boldsymbol{\mathrm{H}}^{1}(\Omega)$-conforming scheme presented in Table 7.1.
@@ Appendix B Divergence-free non-conforming test
Stabilisation using $\mathbb{CR}_{1}-\mathbb{P}_{0},\,\alpha_{0}=20$
971 0.141 4.14e-01 $\star$ 2.24e+00 $\star$ 1.35e+00 $\star$ 3.64e-02 $\star$ 7
3741 0.071 1.99e-01 1.06 1.19e+00 0.91 7.24e-01 0.90 1.82e-02 1.00 7
8311 0.047 1.31e-01 1.04 8.05e-01 0.96 4.91e-01 0.96 1.21e-02 1.00 7
14681 0.035 9.73e-02 1.02 6.08e-01 0.97 3.70e-01 0.98 9.06e-03 1.00 7
22851 0.028 7.76e-02 1.01 4.89e-01 0.98 2.96e-01 0.99 7.24e-03 1.00 7
32821 0.024 6.45e-02 1.01 4.08e-01 0.99 2.47e-01 1.00 6.03e-03 1.00 7
@@
Table B.1: Example B. Error history for a finite element family with $\mathbb{CR}_{1}-\mathbb{P}_{0}-\mathbb{P}_{0}-\mathbb{P}_{1}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})-$norm..

### v2

@@ 8.2 Lagrange stabilisation test
We report on experiments performed using the $\boldsymbol{\mathrm{H}}$-conforming stabilised scheme with Lagrange multipliers proposed in [38] and presented in Section 8.2.
First let us consider the same 2D domain and exact solutions as in Test 8.1 and study the convergence of the conforming scheme using Taylor–Hood elements together with piecewise linear or constant discontinuous elements for the Lagrange multiplier. We also consider stabilised and non-stabilised formulations in order to test the robustness of the scheme. The numerical results portrayed in Tables 8.48.5 clearly confirm the theoretical $O(h^{k+1})$-convergence in the energy norm similarly to the observed/predicted in [39, 37]. The blocks in Table 8.4 show the error history displaying number of degrees of freedom, individual absolute errors, rates of convergence, and Newton iteration counts for the conforming discretisation using Taylor–Hood approximation of velocity-pressure, together with piecewise discontinuous elements for the Lagrange multiplier (on a submesh conforming with the bulk mesh), and piecewise quadratic and continuous functions for the concentration. The choice of $\mathbb{P}_{0}$ for the Lagrange multiplier shows an experimental rate of $\mathcal{O}(h^{1.5})$ for all cases.
On the other hand, the results of using linear discontinuous elements, presented in Table 8.5 show a noticeable deterioration of the convergence for the Lagrange multiplier when the stabilisation is removed. In turn, an optimal rate of convergence $O(h^{k+1})$ is achieved with stabilisation.
We conclude this experiment by showing the result of using the pair $\mathbb{CR}_{1}-\mathbb{P}_{0}$ and $\alpha=0$. Since $\mathbb{CR}_{1}$ is divergence-free, we can use it to approximate (48) without stabilisation. The results obtained for this case are shown in Table 8.6. Here we observe that the convergence for velocity, pressure and concentration is $\mathcal{O}(h)$, while for the Lagrange multiplier we obtain $\mathcal{O}(h^{2})$, similar to that of Tables 8.48.5. It is noteworthy to observe that this scheme is more computationally efficient than Taylor–Hood but not than, e.g., the MINI-element (see Section 8.5 below).
@@
Table 8.4: Example 8.2. Error history for a FE family with $\mathbb{P}^{2}_{2}-\mathbb{P}_{1}-\mathbb{P}_{0}-\mathbb{P}_{2}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm.
@@
88 0.707 7.29e-01 $\star$ 5.21e-01 $\star$ 1.00e+$0$ $\star$ 8.25e-02 $\star$ 6
@@
Table 8.5: Example 8.2. Error history for a FE family with $\mathbb{P}^{2}_{2}-\mathbb{P}_{1}-\mathbb{P}_{1}-\mathbb{P}_{2}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm.
@@
No stabilisation
51 0.707 1.62e+$0$ $\star$ 5.34e-01 $\star$ 7.05e-01 $\star$ 1.81e-01 $\star$ 6
103 0.471 1.13e+$0$ 0.88 3.55e-01 1.00 4.48e-01 1.12 1.22e-01 0.97 7
261 0.283 7.02e-01 0.93 1.85e-01 1.28 2.08e-01 1.50 7.33e-02 1.00 7
793 0.157 3.96e-01 0.97 8.55e-02 1.32 7.36e-02 1.77 4.04e-02 1.01 7
2721 0.083 2.11e-01 0.99 4.02e-02 1.19 2.19e-02 1.90 2.13e-02 1.01 7
10033 0.043 1.09e-01 1.00 1.98e-02 1.07 5.99e-03 1.96 1.10e-02 1.00 7
@@
Table 8.6: Example 8.2. Error history for a FE family with $\mathbb{CR}_{1}-\mathbb{P}_{1}-\mathbb{P}_{0}-\mathbb{P}_{1}$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm. In this case, we take $\delta=1$.
8.3 Filtration with osmotic effects
[…]
With respect to the boundary conditions on the inlet, we consider the following:
$\boldsymbol{u}\cdot\boldsymbol{n}=\left(u_{0}-v_{w}\frac{2x}{d}\right)\left[\frac{3}{2}(1-\lambda^{2})\right]\left[1-\frac{\mathrm{Re}}{420}\left(2-7\lambda^{2}-7\lambda^{4}\right)\right],$
$\boldsymbol{u}_{2,h}$
$\theta_{h}$Refer to caption
To capture the velocity behaviour at the inlet, as well as the maximum permeate velocity, a free tangential stress $(\boldsymbol{\sigma}\boldsymbol{n})\cdot\boldsymbol{t}=0$ is imposed at $\Gamma_{\mathrm{in}}$. Similar results for the dual membrane channel are obtained if we consider $\boldsymbol{u}_{2}$ as the corresponding velocity component on a Berman flow.
The results for the first and second scenarios are depicted in Figure 8.1. For the first case we can see that the velocity near the membrane is affected by the porosity and the transmembrane pressure. Due to the minimal amount of salt compared to the rest of the membrane, at the inlet we observe a high flux in the normal direction to the membrane. In Figure 8.2 we compare the performance of both channels. The concentration profile and permeate velocity are highly dominated by the transmembrane pressure, irrespective of the choice of inlet profiles. On the other hand, we observe that the concentration profile at the membrane increases as we approach the end of the channel, consequently decreasing the permeate velocity. This is accompanied by a linear pressure drop, which behaves similarly for both scenarios.
8.4 A channel with a spacer
We now study the behaviour of the model when an obstacle, serving a a spacer located in the middle of the channel, is considered. The spacer corresponds to the cross section of a cylinder, i.e., a circle, with diameter $3.6\cdot 10^{-4}$m and tangent to the membrane. The boundary conditions for the spacer are the same as $\Gamma_{\mathrm{wall}}$. The velocities to be tested in this experiment are $u_{0}=5\cdot 10^{-2}$m//s and $u_{0}=1.29\cdot 10^{-1}$m//s, and the effect of the salt concentration boundary layer along the membrane is studied.
$|\boldsymbol{u}_{h}|$
$\theta_{h}$Refer to caption
Figure 8.3: Velocity streamlines (top panels) and concentration profiles (bottom panels) around the spacer in a cavity-type configuration and inlet velocities $u_{0}=5.0\cdot 10^{-2}$m//s (left) and $u_{0}=1.29\cdot 10^{-1}$m//s. The bottom numbers indicate distance from inlet (in $\times 10^{-3}$m).
8.5 Benchmark on a 3D geometry
We conclude with a simple benchmark test to compare the effectiveness of the different numerical schemes discussed in the paper. We focus on the unit cube $\Omega:=(0,1)^{3}$ partitioned into $18458$ regular tetrahedra and having $610$ facets on $\Gamma_{\mathrm{m}}$ (similar in size to the fourth mesh level on the 3D experiments in Section 8.1).
We measure the CPU time needed for matrix assembly (denoted Assembly), the percentage of non-zero entries in the matrix (Non-zero pct) and the CPU time to solve the system (Solve). This information is provided by the Python library Time in conjunction with the FEniCS DEBUG log. In the case of the PETSc library, we employ the PETSc Krylov solver for handling the linear system, coupled with the direct solver MUMPS. These metrics are captured during the execution of the first Newton iteration.
For the test, we consider the div-free schemes discussed in Section 4.1 (which we denote by BDM-divfree and CR-divfree), together with the Lagrange multiplier stabilised schemes based on Taylor–Hood elements denoted by TH-stabilised-Pkk (or TH-nonstabilised-Pkk for the $\alpha_{0}=0$ variation), $k=0,1$. The $k$ represents the degree of the polynomial order for the Lagrange multiplier space. Similarly, we denote by MINI-stabilised-Pkk (resp. MINI-nonstabilised-Pkk) the variant where the MINI-element family is used. Finally, the results coming from the first Newton iteration on the scheme (49) are denoted by CR-nonstabilised.
The results are portrayed in Table 8.7. We observe that the divergence-free schemes have a smaller assembly time, but the extra entries (due to the facet jumps and averages) yields a solving time of 16s approximately. Perhaps surprisingly, CR-nonstabilised has the best assembly time of all the test cases, and the solve time is half of its div-free variant. On the other hand, it is important to note the performance hit when considering the stabilisation of (48). The solving time is much higher for the non-stabilised variations (except CR), even failing to converge in the case of MINI-nonstabilised-P1. A clear performance advantage of the MINI family is observed, which is expected because they are the cheapest inf-sup stable FE. The non-zero percentage was kept below $0.1\%$ in all cases, but the TH variants show the lowest sparsity pattern. Also, the div-free and CR-nonstabilised schemes are the ones with the lowest amount of non-zero entries in the global matrix. The results suggest that BDM-divfree, CR-divfree and CR-nonstabilised are preferred if $\nabla\cdot\boldsymbol{u}_{h}=0$ is required, while MINI schemes should be used otherwise. We also tested on the Newton iterations required to solve (49) as a function of $\delta$, where we observed no significant difference. Here we took $\alpha_{0}=0.1$, similarly to Section 8.2.
Table 8.7: Example 8.5. Benchmark of the different methods for solving the coupled flow-transport problem in $\Omega:=(0,1)^{3}$ with $18458$ tetrahedra and $610$ facets on $\Gamma_{\mathrm{m}}$. The symbol – represents a failed Newton iteration.
Scheme DoF Assembly (s) Solve (s) Non-zero pct (%)
BDM-divfree 139391 3.05455 16.6812 0.04734
CR-divfree 139391 2.10487 16.7165 0.04734
CR-nonstabilised 139391 0.87892 7.25941 0.01856
TH-stabilised-P0 118431 5.00461 9.84011 0.08100
TH-stabilised-P1 119651 5.02505 10.2191 0.08005
TH-nonstabilised-P0 118431 3.82103 18.4420 0.08027
TH-nonstabilised-P1 119651 3.80423 25.4627 0.07926
MINI-stabilised-P0 76349 3.18728 2.31272 0.05787
MINI-stabilised-P1 77569 3.13314 2.24265 0.05700
MINI-nonstabilised-P0 76349 2.63304 2.49390 0.05614
MINI-nonstabilised-P1 77569 2.66941 0.05512
CRediT authorship contribution statement
A. Khan, D. Mora, R. Ruíz-Baier, J. Vellojin: Conceptualization, Methodology, Writing-original draft, Review &\& Editing.
Data availability statement
The data that support the findings of this research are available from the corresponding author upon reasonable request.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.